OpenStudy (anonymous):

will give a medal 2(n - 3) + 5 = 3(n - 1)

OpenStudy (anonymous):

so far i have that u distribute 2n-6+5=3n-3

OpenStudy (anonymous):

but what do i do from there

OpenStudy (anonymous):

plz help me

Directrix (directrix):

2n-6+5 = 3n-3 On the left side, what is -6 + 5 = ?

Directrix (directrix):

Answer just what this sum is: -6 + 5 =

Directrix (directrix):

@lubavasilyuk

OpenStudy (anonymous):

why did u flip flop the 6 and 5?

OpenStudy (anonymous):

oh wait sorry

OpenStudy (anonymous):

-6 plus 5 is -1

Directrix (directrix):

2n-6+5 = 3n-3 2n -1 = 3n - 3 The object is to get the n terms on one side and the constant terms on the other 2n -1 = 3n - 3 +1 +1 ---> Add one to both sides and post what you get --------------

OpenStudy (anonymous):

I think the answer is n = 2. Is it right?

Directrix (directrix):

I don't know because I have not worked the problem yet. 2n-6+5=3n-3 2n - 1 = 3n - 3 2n = 3n - 3 + 1 2n = 3n - 2 2n - 3n = -2 next step?

Directrix (directrix):

How do you know that n does not equal -2?

OpenStudy (anonymous):

-1n=-2?

OpenStudy (anonymous):

i get how u are solving everyhing its just that next part the 2n-3n=-2

OpenStudy (anonymous):

hellooooo

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

Okay. 2n-6+5=3n-3 2n - 1 = 3n - 3 2n = 3n - 3 + 1 2n = 3n - 2 2n - 3n = -2 -n = -2 Multiply both sides by (-1) (-1)*(-n) = (-1)*(-2) n = 2

OpenStudy (anonymous):

are u 100 percent sure?

Directrix (directrix):

Let's check it in the original equation: 2(n - 3) + 5 = 3(n - 1) In place of n, put 2. 2(2 - 3) + 5 = 3(2 - 1) 2*(-1) + 5 = 3*(1) ----> Is this statement true or false @lubavasilyuk

Directrix (directrix):

Is -2 + 5 = 3? If so, then n = 2 is correct.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

omg thank u so much

Directrix (directrix):

You are welcome.